Suppose that the USD/ZAR spot rate is quoted as 13.50ZAR and the GBP/USD is $1.30. The cross exchange rate between GBP/ZAR is?
Answer #1

GBP/ZAR = GBP/USD*USD/ZAR = 1.3*13.5=17.55